# Mixup Model Merge: Enhancing Model Merging Performance through Randomized Linear Interpolation

### v1

@@ Abstract
Model merging integrates the parameters of multiple models into a unified model, combining their diverse capabilities. Existing model merging methods are often constrained by fixed parameter merging ratios. In this study, we propose Mixup Model Merge (M3), an innovative approach inspired by the Mixup data augmentation technique. This method merges the parameters of two large language models (LLMs) by randomly generating linear interpolation ratios, allowing for a more flexible and comprehensive exploration of the parameter space. Extensive experiments demonstrate the superiority of our proposed M3 method in merging fine-tuned LLMs: (1) it significantly improves performance across multiple tasks, (2) it enhances LLMs’ out-of-distribution (OOD) robustness and adversarial robustness, (3) it achieves superior results when combined with sparsification techniques such as DARE, and (4) it offers a simple yet efficient solution that does not require additional computational resources. In conclusion, M3 is a simple yet effective model merging method that significantly enhances the performance of the merged model by randomly generating contribution ratios for two fine-tuned LLMs. The code is available at https://github.com/MLGroupJLU/MixupModelMerge.
@@ 1 Introduction
(c) The performance trend of the merged models obtained through the following methods: TIES-Merging without M3 and DARE, TIES-Merging with DARE, TIES-Merging with M3, and TIES-Merging with both M3 and DARE.
Figure 1: (a) M3 significantly boosts the model merging performance. (b) OOD robustness, and adversarial robustness of the merged models. (c) Combined with DARE, M3 delivers even better results. LM, MATH, and Code refer to the WizardLM13B, WizardMath-13B, and llama-2-13b-code-alpaca models. TA stands for Task Arithmetic and TIES stands for TIES-Merging.
@@
Supervised Fine-Tuning (SFT) is a crucial technique for adapting LLMs to specific tasks, refining their performance by training on domain-specific data (Hu et al., 2021; Ding et al., 2023; Xia et al., 2024). However, SFT requires substantial computational resources and long training times (Brown et al., 2020; Chang et al., 2024a). To address this challenge, Model Merging has emerged as an efficient solution, fusing the parameters of multiple fine-tuned LLMs into a unified model with diverse capabilities, without the need for additional training or computational costs (Yang et al., 2024; Akiba et al., 2024). It effectively reduces the resource-intensive demands of SFT while preserving and even enhancing model performance.
A simple analogy for model merging is the Super Mario game, where the protagonist gains special abilities by absorbing power-up items. Similarly, merging model parameters integrates the strengths of different models, enabling more effective multi-task learning (Yu et al., 2024). However, existing model merging methods have some limitations, these methods heavily rely on predefined or heuristic parameter fusion strategies (Wortsman et al., 2022; Ilharco et al., 2022; Matena and Raffel, 2022; Jin et al., 2022; Yadav et al., 2023; Yu et al., 2024) such that they fail to fully explore the parameter space, thereby restricting the potential of the merged model.
@@
To address this issue, we draw inspiration from Mixup (Zhang, 2017) and propose a novel technique called Mixup Model Merge (M3). This method introduces randomness by dynamically adjusting the contribution ratios between models, making the model merging process more flexible and enabling a thorough exploration of the parameter space. M3 further unlocks the potential of model merging, significantly enhancing the generalization performance of the merged model while improving its out-of-distribution (OOD) and adversarial robustness (Wang et al., 2023; Zhu et al., 2023). As shown in Figure 2, the implementation of M3 is similar to the process of proportionally mixing magical potions in Harry Potter. Specifically, this method dynamically controls the parameter fusion ratio between two fine-tuned LLMs by randomly generating a linear interpolation ratio $\lambda_{m}$, where $\lambda_{m}\in(0,1)$ and $\lambda_{m}\sim\text{Beta}(\alpha,\alpha)$. By adjusting the parameter $\alpha$, we can precisely control the distribution of $\lambda_{m}$, allowing M3 to explore the parameter space of model merging more deeply and thus fully unleash the potential of the models, leading to improved overall performance.
@@
We conducted extensive experiments with three fine-tuned LLMs: WizardLM-13B (Xu et al., 2024), WizardMath-13B (Luo et al., 2023), and llama-2-13b-code-alpaca (Chaudhary, 2023), which specialize in instruction following, mathematical reasoning, and code generation, respectively. Inspired by Mixup’s effectiveness in enhancing the robustness of neural networks when handling corrupted labels or adversarial examples (Zhang, 2017), we further performed comprehensive evaluations on LiveBench (White et al., 2024) and PromptBench (Zhu et al., 2024) to validate the potential of M3 in improving the OOD robustness and adversarial robustness of merged models. The experimental results demonstrate that our proposed M3 method can significantly improve merged models’ performance across various tasks (as shown in Figure 1(a)), enhance the OOD and adversarial robustness of the merged models (as shown in Figure 1(b)), and boost model merging when combined with sparsification techniques like DARE (Yu et al., 2024) (as shown in Figure 1(c)).
@@ 2 Related Works
Model merging is a technique that integrates the parameters of multiple models to create a unified model with enhanced or diverse capabilities (Wortsman et al., 2022; Ilharco et al., 2022; Matena and Raffel, 2022; Jin et al., 2022; Yadav et al., 2023; Yu et al., 2024; Lin et al., 2024). Task arithmetic (Ilharco et al., 2022) leverages task vectors for model merging through arithmetic operations, incorporating a predefined scaling term to weight the contribution of different models. Fisher Merging (Matena and Raffel, 2022) performs parameter fusion by applying weights derived from the Fisher information matrix (Fisher, 1922), resulting in more precise parameter integration. TIES-Merging (Yadav et al., 2023) addresses task conflicts by removing low-magnitude parameters, resolving sign disagreements, and merging only the parameters that align with the final agreed-upon sign. In (Yu et al., 2024), it is found that LLMs can enhance their capabilities through model merging. Additionally, it introduces DARE, a method for sparsifying the delta parameters of the model (Ilharco et al., 2022), significantly improving the performance of various model merging techniques.
Mixup
Mixup is proposed to enhance the generalization ability of deep learning models by surpassing traditional Empirical Risk Minimization (ERM) (Zhang, 2017). It is a simple, data-agnostic augmentation technique that trains models using virtual examples created by linearly interpolating pairs of random examples and their corresponding labels. Rooted in the Vicinal Risk Minimization (VRM) principle (Chapelle et al., 2000), this approach improves generalization across a variety of datasets (Russakovsky et al., 2015; Krizhevsky et al., 2009; Warden, 2017; Asuncion et al., 2007), and helps reduce overfitting, sensitivity to adversarial examples, and training instability, all with minimal computational cost. Given two samples $(x_{i},y_{i})$ and $(x_{j},y_{j})$, Mixup generates a new sample using the following formulas:
[…]
$\tilde{x}$ and $\tilde{y}$ represent the generated synthetic sample and its label, respectively, with $\lambda$ determining their interpolation ratio, typically ranging from 0 to 1. Here, $\lambda$ is a hyperparameter sampled from a Beta distribution, i.e., $\lambda\sim\text{Beta}(\alpha,\alpha)$, where $\alpha$ controls the strength of the interpolation between feature-target pairs. Inspired by Mixup, we propose a novel model merging method, M3, which generates the parameters of the merged model by performing linear interpolation between the parameters of two fine-tuned LLMs, with the interpolation ratio being random.
@@ 3 Methodology
3.1 Model Merging Problem
Following Yu et al. (2024), we focus on merging fine-tuned LLMs that have been optimized from the same pre-trained backbone (Touvron et al., 2023). Specifically, we aim to fuse the parameters of these LLMs to create a unified model capable of handling multiple tasks. In this context, we restrict our attention to the merging of two models, as the mixup theory, which forms the basis of our approach, is generally applied to two entities.
Given two tasks $t_{1}$ and $t_{2}$ with the corresponding fine-tuned LLMs having parameters $\theta_{\text{SFT}}^{t_{1}}$ and $\theta_{\text{SFT}}^{t_{2}}$, model merging aims to combine the parameters of these two models into a single model with parameters $\theta_{M}$. The resulting model should be able to effectively perform both tasks simultaneously, leveraging the knowledge learned from each individual model.
@@ 3.2 Mixup Model Merge
Inspired by Mixup (Zhang, 2017), we propose a simple yet effective model merging method called Mixup Model Merge (M3). Unlike existing methods that use fixed merging ratios, M3 generates the model contribution ratio randomly, harnessing randomness to inject fresh vitality into the model merging process. Specifically, M3 further explores the parameter space of the merged model to unlock the potential of model merging.
To further elaborate, given two fine-tuned LLMs with parameters $\{\theta_{\text{SFT}}^{t_{1}},\theta_{\text{SFT}}^{t_{2}}\}$, we combine M3 with established model merging methods to fuse these parameters and obtain a single merged model with parameters $\theta_{M}$. As illustrative examples, we consider two widely used merging methods: Average Merging (Wortsman et al., 2022) and Task Arithmetic (Ilharco et al., 2022).
@@
where $\theta_{\text{PRE}}\in\mathbb{R}^{d}$ represents the parameters of the pre-trained language model (PLM), such as Llama 2 (Touvron et al., 2023). $\lambda$ is a scaling factor that weights the contribution of each model during the merging process. $\delta^{t_{i}}$ denotes the delta parameter (Ilharco et al., 2022), which is defined as the difference between the parameters of the language models (LMs) before and after SFT, i.e., $\delta^{t}=\theta_{\text{SFT}}^{t}-\theta_{\text{PRE}}\in\mathbb{R}^{d}$, where $t$ refers to task $t$.
@@
where $\lambda_{m}$ determines the linear interpolation ratio between the two fine-tuned LLMs, and is generally a value between 0 and 1. $\lambda_{m}$ is sampled from a Beta distribution, typically $\lambda_{m}\sim\text{Beta}(\alpha,\alpha)$, where $\alpha$ controls the shape of the Beta distribution.
The hyperparameter $\alpha$ for M3 is selected from the range $[0.2,0.4,0.5,1,2,3,5]$. As shown in Figure 3: (1) When $\alpha=1$, $\lambda$ follows a uniform distribution, meaning all values within the range $(0,1)$ are equally likely to be sampled. (2) When $\alpha<1$, the distribution of $\lambda$ exhibits a bimodal shape, with higher probabilities near the extremes (0 and 1). This indicates that the merged model is more likely to be dominated by one of the two models. (3) When $\alpha>1$, the distribution of $\lambda$ becomes concentrated around the middle (e.g., 0.5), resulting in more balanced contributions from both models.
@@ 3.3 Theoretical Analysis
Mixup performs linear interpolations between data samples and their labels in the data space (Zhang, 2017). Similarly, M3 can be viewed as applying random linear interpolation in the parameter space, where the interpolation occurs between the parameters of two fine-tuned models trained on different tasks. M3 represents a natural extension of the Vicinal Risk Minimization (VRM) principle (Chapelle et al., 2000) into the model parameter space. In data space, VRM introduces a vicinal distribution to simulate the true data distribution, thereby increasing the diversity of training data and enhancing the model’s generalization ability. Similarly, M3 extends this principle by constructing a virtual neighborhood in the model parameter space between two task-specific models, combining their knowledge in a way that is more natural and balanced.
In the context of model merging, interpolating between two sets of model parameters, $\theta_{\text{SFT}}^{t_{1}}$and $\theta_{\text{SFT}}^{t_{2}}$, defines a new neighborhood distribution in the parameter space, which can be expressed as:
[…]
where $P_{\nu}(\theta_{M})$ represents the probability distribution of the new model parameters $\theta_{M}$ generated through interpolation. The function $\nu(\theta_{M}\mid\theta_{\text{SFT}}^{t_{1}},\theta_{\text{SFT}}^{t_{2}})$ defines the range and behavior of $\theta_{M}$, depending on the interpolation strategy (such as linear interpolation). Additionally, $P(\theta_{\text{SFT}}^{t_{1}})$ and $P(\theta_{\text{SFT}}^{t_{2}})$ represent the prior distributions of the parameters of the two models, respectively.
Under the linear interpolation strategy, the function $\nu(\theta_{M}\mid\theta_{\text{SFT}}^{t_{1}},\theta_{\text{SFT}}^{t_{2}})$ is defined as:
$\theta_{M}=\lambda_{m}\cdot\theta_{\text{SFT}}^{t_{1}}+(1-\lambda_{m})\cdot\theta_{\text{SFT}}^{t_{2}},$ (7)
where $\lambda_{m}\in(0,1)$ is the interpolation ratio. The corresponding distribution $P_{\nu}(\theta_{M})$ is expressed as:
where $\delta(\cdot)$ is the Dirac delta function, ensuring that $\theta_{M}$ satisfies the linear interpolation rule. By using this linear interpolation method, the parameters $\theta_{\text{SFT}}^{t_{1}}$ and $\theta_{\text{SFT}}^{t_{2}}$ are combined in the parameter space, forming a new neighborhood distribution $P_{\nu}(\theta_{M})$. This process effectively merges the knowledge of both models into a unified parameter space, thus achieving robust performance across different tasks while maintaining the original strengths of the models.
By interpolating the parameters, M3 encourages the model to learn smoother decision boundaries. In this context, smoother decision boundaries can be understood as the boundaries between different tasks, such as instruction following, mathematical reasoning, and code generation, where the model must understand and adapt to each task differently. In tasks $t_{1}$ and $t_{2}$, M3 creates a virtual neighborhood that seamlessly integrates knowledge from both tasks. This approach prevents the merged model from overfitting task-specific details, ensuring a balanced and effective performance across all tasks.
Secondly, M3 introduces a linear inductive bias in the parameter space, encouraging the merged model parameters to lie on a linear manifold between the two source models. This linear structure offers significant advantages in terms of simplicity and generalization. According to Occam’s Razor, simpler solutions tend to generalize better. By performing linear interpolation between two sets of parameters, M3 avoids unnecessary complexity in the model merging process, leading to a more straightforward and efficient solution.
Thirdly, M3 can improve the performance of the merged model across multiple tasks by mitigating task conflicts. Different tasks may require conflicting parameter values, which can lead to performance degradation on one task while optimizing for another. Linear interpolation helps balance these conflicts, resulting in a model that performs well among all tasks.

### v2

@@ Abstract
Model merging aims to integrate multiple task-specific models into a unified model that inherits the capabilities of the task-specific models, without additional training. Existing model merging methods often lack consideration of the varying contribution ratios of different task-specific models to the final merged model. In this paper, we propose Mixup Model Merge (M3), a simple yet effective method inspired by the randomized linear interpolation strategy from the Mixup data augmentation technique. M3 performs randomized linear interpolation in parameter space between two task-specific LLMs, where interpolation coefficients are sampled from a Beta distribution to explore diverse contribution ratios. This controllable randomness allows M3 to outperform standard equal-ratio merging by discovering better contribution ratio combinations. Extensive experiments show that M3 significantly (1) improves merged LLM performance across tasks, (2) enhances out-of-distribution and adversarial robustness, and (3) outperforms the positive effects of the sparsification method DARE on model merging and can be further combined with DARE to achieve superior results. By tuning the Beta distribution’s shape parameters, (4) M3 balances exploration efficiency and diversity in contribution ratios. The code is available at https://github.com/MLGroupJLU/MixupModelMerge.
@@ 1 Introduction
Figure 1: (a) M3 significantly boosts the model merging performance. (b) OOD robustness, and adversarial robustness of the merged models.
@@
Supervised Fine-Tuning (SFT) is a crucial technique for adapting LLMs to specific tasks, refining their performance by training on domain-specific data (Hu et al., 2021; Ding et al., 2023; Xia et al., 2024). However, SFT requires substantial computational resources and long training times (Brown et al., 2020; Chang et al., 2024a). To address this challenge, Model Merging has emerged as an efficient solution, fusing the parameters of multiple fine-tuned LLMs into a unified model with diverse capabilities, without the need for additional training or computational costs (Yang et al., 2024; Akiba et al., 2024). However, Existing model merging methods often adopt identical contribution ratios for different task-specific models in the merging process (Wortsman et al., 2022; Ilharco et al., 2022; Matena and Raffel, 2022; Jin et al., 2022; Yadav et al., 2023; Yu et al., 2024), or rely on a few manually predefined contribution ratio values (Lin et al., 2023). This restricts the exploration of how varying contribution ratios influence the final merged model, and limits the ability to fully explore the resulting parameter space. Consequently, the potential of the merged model is significantly constrained.
@@
To address the above issue, we propose a novel approach, Mixup Model Merge (M3), for exploring appropriate contribution ratios among task-specific models. This method is inspired by the randomized linear interpolation strategy from the Mixup data augmentation technique (Zhang, 2017). In contrast to conventional Mixup, M3 performs linear interpolation in the model parameter space, where the interpolation coefficients are randomly sampled from a Beta distribution. Specifically, this method controls the contribution ratio between two fine-tuned LLMs by randomly generating a linear interpolation ratio $\lambda_{m}$, where $\lambda_{m}\in(0,1)$ and $\lambda_{m}\sim\text{Beta}(\alpha,\alpha)$. By adjusting the parameter $\alpha$, we can precisely control the distribution of $\lambda_{m}$, enabling M3 to balance the diversity of $\lambda_{m}$ values with the efficiency of discovering suitable contribution ratios. In the parameter space, our method explores the family of models formed along the linear path between two task-specific LLMs by randomly sampling interpolation coefficients. This allows us to identify suitable interpolation coefficients (i.e., contribution ratios), thereby obtaining a merged model with superior performance. As shown in Figure 2, the implementation of M3 is similar to the process of proportionally mixing magical potions in Harry Potter.
@@
We conducted extensive experiments with three homologous task-specific fine-tuned LLMs: WizardLM-13B (Xu et al., 2024), WizardMath-13B (Luo et al., 2023), and llama-2-13b-code-alpaca (Chaudhary, 2023), which specialize in instruction following, mathematical reasoning, and code generation, respectively. Inspired by Mixup’s effectiveness in enhancing the robustness of neural networks when handling corrupted labels or adversarial examples (Zhang, 2017), we further performed comprehensive evaluations on LiveBench (White et al., 2024) and PromptBench (Zhu et al., 2024) to validate the potential of M3 in improving the OOD robustness and adversarial robustness (Wang et al., 2023; Zhu et al., 2023) of merged models. The experimental results demonstrate that our proposed M3 method can significantly improve merged models’ performance across various tasks (as shown in Figure 1(a)), and enhance the OOD and adversarial robustness of the merged models (as shown in Figure 1(b)).
@@ 2 Related Works
Model merging is a technique that integrates the parameters of multiple models to create a unified model with enhanced or diverse capabilities (Wortsman et al., 2022; Ilharco et al., 2022; Matena and Raffel, 2022; Jin et al., 2022; Yadav et al., 2023; Yu et al., 2024; Lin et al., 2024). Task arithmetic (Ilharco et al., 2022) leverages task vectors for model merging through arithmetic operations, incorporating a predefined scaling term to weight the contribution of different models. Fisher Merging (Matena and Raffel, 2022) performs parameter fusion by applying weights derived from the Fisher information matrix (Fisher, 1922), resulting in more precise parameter integration. TIES-Merging (Yadav et al., 2023) addresses task conflicts by removing low-magnitude parameters, resolving sign disagreements, and merging only the parameters that align with the final agreed-upon sign. In (Yu et al., 2024), it is found that LLMs can enhance their capabilities through model merging. Additionally, it introduces DARE, a method for sparsifying the delta parameters of the model (Ilharco et al., 2022), significantly improving the performance of various model merging techniques. DELLA (Deep et al., 2024) is a novel model merging technique that integrates a new pruning strategy called MAGPRUNE, which samples delta parameters based on their magnitudes. MAGPRUNE demonstrates improvements over existing methods such as DARE and TIES.
@@ 3 Methodology
We focus on merging the parameters of two homologous task-specific fine-tuned LLMs to construct a unified model that effectively preserves the capabilities of both, without requiring any further training.
@@ 3.1 Mixup Model Merge
Inspired by the random interpolation strategy of Mixup (Zhang, 2017), we extend the idea from input space to parameter space, proposing Mixup Model Merge (M3).
We consider two task-specific fine-tuned language models, denoted as $\theta_{t_{1}}^{\text{SFT}}$ and $\theta_{t_{2}}^{\text{SFT}}$, obtained by independently fine-tuning a shared pretrained base model $\theta_{\text{PRE}}$ on two different tasks $t_{1}$ and $t_{2}$, respectively. To explore the continuum of models between these two experts, we define a linear interpolation over their parameters:
$\theta_{M}=\lambda_{m}\theta_{\text{SFT}}^{t_{1}}+(1-\lambda_{m})\theta_{\text{SFT}}^{t_{2}},\quad\lambda_{m}\in(0,1)$
This defines a one-dimensional family of interpolated models. Alternatively, the interpolation can be expressed in the delta space, where we defined as the parameter offsets between the task-specific fine-tuned model and the pretrained model:
$\delta_{t_{i}}=\theta_{\text{SFT}}^{t_{i}}-\theta_{\text{PRE}},\quad i\in\{1,2\}$
Then the interpolated model can be reformulated as:
$\theta_{M}=\theta_{\text{PRE}}+\lambda_{m}\delta_{t_{1}}+(1-\lambda_{m})\delta_{t_{2}}$
In both formulations, the resulting model $\theta_{M}$ lies on the line segment connecting the two task-specific models in parameter space. This construction naturally defines a one-dimensional model family:
$\mathcal{F}_{M}=\{\theta_{M}^{\lambda_{m}}\mid\lambda_{m}\in(0,1)\}$
Such a model family allows us to study how task behaviors interact and mix at the parameter level, and serves as the foundation for our proposed Mixup Model Merge ($M^{3}$).
Linear Interpolation Coefficient
A key component of the M3 method is the stochastic sampling of the interpolation coefficient $\lambda_{m}$ from a Beta distribution, i.e., $\lambda_{m}\sim\text{Beta}(\alpha,\alpha)$. Defined over the interval $(0,1)$, the Beta distribution is naturally suited for interpolation and introduces controllable randomness into the merging process. Its tunable shape parameter $\alpha$ allows us to adjust the sampling bias of $\lambda_{m}$, enabling targeted and efficient exploration of the one-dimensional model space $\mathcal{F}_{M}$ for optimal model combinations. Specifically, as illustrated in Figure 3:
  • When $\alpha<1$, the distribution of $\lambda_{m}$ exhibits a bimodal shape, with higher probabilities near the extremes (0 and 1), indicating that the merged model is more likely to be dominated by one of the two task-specific models. As $\alpha$ decreases, the range of sampled $\lambda_{m}$ values becomes more concentrated, leading to higher exploration efficiency but lower diversity of the merged models explored.
  • When $\alpha>1$, the distribution of $\lambda_{m}$ is concentrated around the middle region (e.g., near 0.5), resulting in more balanced contributions from both models. As $\alpha$ increases, the range of sampled $\lambda_{m}$ values also becomes more concentrated, improving exploration efficiency while reducing the diversity of merged models.
  • When $\alpha=1$, $\lambda_{m}$ follows a uniform distribution, meaning that all values within the interval $(0,1)$ are equally likely to be sampled. As $\alpha$ approaches 1, the distribution of $\lambda_{m}$ becomes more dispersed, leading to lower exploration efficiency but higher diversity in the merged models explored.
Comparison with Grid Search
As shown in Table 1, we compare the performance of M3 with that of grid search. Grid search requires manually specifying multiple fixed and discrete interpolation coefficients (e.g., $\lambda=0.1,0.2,\ldots,0.9$). While this approach offers fine-grained control, it lacks the ability to explore continuous values. Achieving finer resolution (e.g., identifying values beyond the first decimal place) requires a smaller step size, which significantly reduces search efficiency.
In contrast, our method enables more efficient and diverse exploration of continuous interpolation coefficients. For example, our approach can sample values such as $\lambda_{m}=0.62156685$, which are difficult to obtain with coarse grid search. Empirically, we find that the model’s performance is often sensitive to the second decimal place of $\lambda$. To achieve such granularity, grid search would require a step size of 0.01, resulting in a prohibitively large number of evaluations.
Our method, by comparison, typically requires only a single pass over a small set of hyperparameter values for $\alpha$ (e.g., $\alpha\in\{0.2,0.4,0.5,1,2,3,5\}$). In cases where a suitable $\lambda_{m}$ is not found in the first pass, we perform a few additional samplings under a selected $\alpha$ value—no more than three times in our experiments. In Table 2, we report the specific $\lambda_{m}$ values that lead to performance improvements, along with the corresponding distribution shape parameter $\alpha$ and the total number of sampling iterations.
Method Distribution of $\lambda_{m}$ Continuity of $\lambda_{m}$ Efficiency Diversity Controllability
Grid Search Fixed discrete values (e.g., 0.1–0.9) No Low Medium High
M3+Beta(1,1) Uniform over (0,1) Yes Medium High Medium
M3+Beta(5,5) Concentrated near the center (0.5) Yes High Low High
M3+Beta(0.2,0.2) Concentrated near the endpoints (0 or 1) Yes High Low High
Table 1: Comparison of interpolation coefficient $\lambda_{m}$ sampling strategies.
M3 Implementation
M3 is a plug-and-play approach that can be integrated into existing representative model merging methods to enhance the performance of the resulting merged models. As illustrative examples, we consider two widely used merging methods: Average Merging (Wortsman et al., 2022) and Task Arithmetic (Ilharco et al., 2022).
@@
$\lambda$ is a scaling term that takes values from the set $\{0.0,0.1,\ldots,1.0\}$.
@@
where
$\lambda_{m}\in[0,1],\quad\lambda_{m}\sim\mathrm{Beta}(\alpha,\alpha),$
$\lambda_{m}$ controls the interpolation ratio, i.e., the contribution ratio, between the two fine-tuned models. $\alpha$ determines the shape of the Beta distribution.
@@ 3.2 Theoretical Analysis
According to the theoretical and empirical findings of Neyshabur et al. (2020), fine-tuning models from the same pretrained initialization can guide them into a shared, flat, and connected low-loss basin. As a result, the linear interpolation path between such fine-tuned models remains within this low-loss basin and does not encounter significant performance barriers. In some cases, models along the interpolation path even outperform the endpoints.
Since both $\theta_{t1}^{\mathrm{SFT}}$ and $\theta_{t2}^{\mathrm{SFT}}$ are fine-tuned from the same pretrained model, they can be considered to lie within the same flat and connected low-loss basin. Therefore, linearly interpolating between them is unlikely to cause a collapse in performance, and may even yield interpolated models with better performance than either endpoint. This provides a theoretical justification for exploring interpolated models along the linear path between two fine-tuned models as a means of finding superior merged models.
Consider the $i$-th parameter of the interpolated model:
$\theta_{M,i}=\theta_{\mathrm{PRE},i}+\lambda_{m}\,\delta_{t1,i}+(1-\lambda_{m})\,\delta_{t2,i},$ (5)
where $\delta_{t1,i}=\theta_{t1,i}^{\mathrm{SFT}}-\theta_{\mathrm{PRE},i}$ and $\delta_{t2,i}=\theta_{t2,i}^{\mathrm{SFT}}-\theta_{\mathrm{PRE},i}$ denote the delta parameters from the pretrained model $\theta_{\mathrm{PRE}}$ to the task-specific fine-tuned models $\theta_{t1}^{\mathrm{SFT}}$ and $\theta_{t2}^{\mathrm{SFT}}$, respectively.
Define a function representing the change of the merged model’s $i$-th parameter relative to the pretrained model:
$f_{i}(\lambda_{m})=\lambda_{m}\,\delta_{t1,i}+(1-\lambda_{m})\,\delta_{t2,i}.$ (6)
If $\delta_{t1,i}>0$ and $\delta_{t2,i}<0$, the delta parameters are in conflict and may partially cancel each other. In particular, when
$\lambda_{m}=\frac{|\delta_{t2,i}|}{|\delta_{t1,i}|+|\delta_{t2,i}|}\in(0,1),$ (7)
we have $f_{i}(\lambda_{m})=0$, meaning the conflicting delta parameters are exactly canceled out at that parameter.
Even when $\lambda_{m}$ is close to this cancellation point, the magnitude of the update is significantly reduced. This shows that linear interpolation can effectively suppress conflicting delta parameters at certain parameters, providing stability in the merged model.